# shapes.geometric, shapes.multipart, arrows, arrows.meta, positioning, calc, chains, backgrounds Towards Explainable Evaluation Metrics for Machine Translation

### v1

@@ 3 Literature Review & Selection
[…]
[category] (-0.100000-5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000) rectangle node (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000);
[…]
[title] (-0.100000-4.300000/1.75, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000) rectangle nodequality
aspects (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
[title] (-0.100000-4.300000/1.75, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000) rectangle nodefeature
importance (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
[title] (-0.100000-4.300000/1.75, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000) rectangle nodefine-grained
errors (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000);
[category] (-0.100000-5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000) rectangle node (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000);
[…]
[title] (-0.100000-4.300000/1.75, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000) rectangle nodeperturbation
robustness (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
[title] (-0.100000-4.300000/1.75, 0.100000+ 0.985000+ 0.100000) rectangle nodelinguistic
properties (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000);
[category] (0.100000- 5.100000/2, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000+ 0.100000+ 0.300000) rectangle node post-hoc (0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000+ 0.100000);
[…]
[title] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000) rectangle nodeblack-box (0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000);
[title] (0.100000+ 5.100000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000) rectangle nodegradient (0.100000+ 5.100000+ 0.100000+ 2.800000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000);
[title] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000) rectangle nodeattention
weights (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000);
[title] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000) rectangle nodeembeddings (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000);
[category] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000- 5.100000/2, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000+ 0.100000+ 0.300000) rectangle node intrinsic (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000+ 0.100000);
[…]
[title] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000) rectangle nodemodel specific (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000);
[content] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+0.985000) rectangle node(Thompson and Post 2020)
(Yuan et al. 2021)
(Kamal Eddine et al. 2023)
(Fu et al. 2023) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
[content] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000) rectangle node(Kaster et al. 2021) (0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000);
[content] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000) rectangle node(Golovneva et al. 2023)
(Opitz and Frank 2022) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000);
[content] (0.100000+ 5.100000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000) rectangle node(Rei et al. 2022b)
(Rei et al. 2023) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000);
[content] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000) rectangle node(Treviso et al. 2021) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000);
[content] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+0.985000) rectangle node(Fernandes et al. 2022b)
(Kabir and Carpuat 2021)
(Rubino et al. 2021)
(Wang et al. 2022) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
[content] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+0.985000) rectangle node(Leiter 2021)
(Tao et al. 2022)
(Azadi et al. 2022) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
[content] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000) rectangle node(Eksi et al. 2021) (0.100000+ 5.100000+ 0.100000+ 2.800000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000);
[content] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000) rectangle node(Fomicheva et al. 2022) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000);
[content] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000) rectangle node(Leiter et al. 2022b) (0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000);
[content] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000) rectangle node(Lu et al. 2022, 2023)
(Xu et al. 2023), … (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000);
[content] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+0.985000) rectangle node(Sai et al. 2022)
(Karpinska et al. 2022)
(He et al. 2022)
(Sun et al. 2022)
(Chen and Eger 2023)
… (0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
[content] (0.100000, 0.100000+ 0.985000+ 0.100000) rectangle node(Kaster et al. 2021) (0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000);
[content] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000, 0.100000+ 0.985000+ 0.100000) rectangle node(Opitz and Frank 2022) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000);
[-] (0.100000- 0.100000- 5.100000/1.75 + 0.3, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000) – (0.100000- 0.100000- 5.100000/1.75 - 0.100000+0.3, 0.100000+ 0.985000+ 0.100000) node[anchor=mid, below, rotate=-90, xshift=-6cm] decision-understanding model-understanding;
[set/.style = draw]
[…]
Adequacy: 0.7
Fluency: 1;
[…]
“Things became complicated”
“Die Dinge bekamen kompliziert”;
[…]
“Things sensebecame/sense complicated”
“Die Dinge sensebekamen/sense kompliziert”;
[…]
Numeric: “3 Dinge bekamen kompliziert”lower score
Order: “kompliziert die Dinge bekamen”lower score
[…]
(Example)[set, align=left, rectangle, dotted] at (-6.5, 5.8) Example Data:
Ground Truth: “Things became complicated”
Hypothesis: “Die Dinge bekamen kompliziert”
Score: ;
[…]
@@ 4 Taxonomy
Quality aspects (Row 1 in Figure 4) This category comprises works that present explanations of the metric output by producing separate scores for different aspects of the translation. Most methods that we list develop generation-based metrics that are explainable by design (“model specific” in Figure 4). This means, they do not use an external explainer that could be applied to other metrics. BARTScore (Yuan et al. 2021) uses the autoregressive generation model BART (Lewis et al. 2020) to predict the average word generation probability of a sentence , given a sentence :
This formula sums over the log-probability of token being generated by BART with parameters given all previous tokens of and all tokens of . Depending on whether A is the ground truth sentence and B the hypothesis (or vice-versa), Yuan et al. (2021) state that different aspects of the translation are evaluated. For example, they associate the generation direction of A:source B:hypothesis with the properties coherence and fluency. This idea of using the generation probabilities in multiple directions was first introduced in PRISM (Thompson and Post 2020), but Yuan et al. (2021) formalized the connection of the directions to known quality aspects. DATScore (Kamal Eddine et al. 2023) extends BARTScore by using a multilingual model and further generation directions (additionally using the source sentence). The metric GPTScore (Fu et al. 2023) provides a suite of prompts for language models that emphasize specific aspects when requesting a translation or paraphrase. Then, similar to BARTScore, the generation probabilities are averaged for each word. For example, they use the prompt “Rewrite the following text to make it more grammatical and well-written: [reference or hypothesis] In other words,” to generate a paraphrase whose probabilities they use for a fluency score. The example specifies reference/hypothesis to describe that the metric checks either the generation probability of the hypothesis given the reference, or vice-versa.
@@
Besides adding explanatory value per se, quality aspect explanations have two more potential use cases for explainable metrics: (1) separate quality scores could be aggregated, to compute single, decomposable sentence-level scores and (2) scores could be aggregated over data sets to check which aspects a metric favors the most. For example, Kaster et al. (2021); Opitz and Frank (2022) also assign quality aspect scores, but mostly use them for model understanding; see linguistic properties explanations for more details.
@@
Goals: Quality aspect approaches could be used for metric accessibility and semi-automatic labeling. The usage for metric diagnosis is limited, as the approaches are often based on models that are explainable by design and might not be faithful (the models generate their own explanations). The current approaches cannot be used to detect social biases.
Audiences: The quality aspects fluency and adequacy follow simple definitions, such that these metrics could be used by non-expert and expert users.
Limitations: The generation directions might not fully match the quality aspects. Also, the usage of prompting leaves it to another black box setting whether the model picks up the desired properties or not. Lastly, the explanatory value of quality aspect approaches per se is rather small, as they only allow to reason, “the rated translation is good because of aspect x” and to guess about the model’s reasoning. As none of the current works101010Kaster et al. (2021); Opitz and Frank (2022) are considered at a later point. specifically targets explainability, they do not consider faithfulness. Plausibility is automatically considered in standard evaluation schemes with human judgments.
@@
Feature importance (Row 2 in Figure 4) Feature importance techniques assign scores to all features in the input of a metric (here the input tokens) that express their importance to the output of the metric.111111Other names include relevance scores or attribution scores. Sometimes the scores are used to highlight important words. Fomicheva et al. (2022) note that high feature importance scores should correspond to translation errors, as humans pay most attention to errors when rating a translation. They build on the idea that when humans evaluate translations, they often focus on the errors that they can identify on a word- or phrase-level (Freitag et al. 2021a). Then, Fomicheva et al. (2022) evaluate how well feature importance is correlated with human word-level error annotations, i.e., the plausibility of the explanations.

### v2

@@ 4 Taxonomy
Quality aspects (Row 1 in Figure 3) This category comprises works that present explanations of the metric output by producing separate scores for different aspects of the translation. Most methods that we list develop generation-based metrics that are explainable by design (“model specific” in Figure 3). This means, they do not use an external explainer that could be applied to other metrics. BARTScore (Yuan et al., 2021) uses the autoregressive generation model BART (Lewis et al., 2020) to predict the average word generation probability of a sentence , given a sentence :
This formula sums over the log-probability of token being generated by BART with parameters given all previous tokens of and all tokens of . Depending on whether A is the ground truth sentence and B the hypothesis (or vice-versa), Yuan et al. (2021) state that different aspects of the translation are evaluated. For example, they associate the generation direction of A:source B:hypothesis with the properties coherence and fluency. This idea of using the generation probabilities in multiple directions was first introduced in PRISM (Thompson and Post, 2020), but Yuan et al. (2021)
formalized the connection of the directions to known quality aspects. DATScore (Kamal Eddine et al., 2023) extends BARTScore by using a multilingual model and further generation directions (additionally using the source sentence).
The metric GPTScore (Fu et al., 2023) provides a suite of prompts for language models that emphasize specific aspects when requesting a translation or paraphrase. Then, similar to BARTScore, the generation probabilities are averaged over all tokens. For example, they use the prompt “Rewrite the following text to make it more grammatical and well-written: [reference or hypothesis] In other words,”
to generate a paraphrase whose probabilities they use for a fluency score. The example specifies reference/hypothesis to describe that the metric checks either the generation probability of the hypothesis given the reference, or vice-versa.
@@
Besides adding explanatory value per se, quality aspect explanations have two more potential use cases for explainable metrics: (1) separate quality scores could be aggregated, to compute single, decomposable sentence-level scores and (2) scores could be aggregated over data sets to check which aspects a metric favors the most. For example, Kaster et al. (2021) and Opitz and Frank (2022) also assign quality aspect scores, but mostly use them for model understanding; see linguistic properties explanations for more details.
@@
Feature importance (Row 2 in Figure 3) Feature importance techniques assign scores to all features in the input of a metric (here the input tokens) that express their importance to the output of the metric.101010Other names include relevance scores or attribution scores. Fomicheva et al. (2022) note that high feature importance scores should correspond to translation errors, as humans pay most attention to errors when rating a translation. They build on the idea that when humans evaluate translations, they often focus on the errors that they can identify on a word- or phrase-level (Freitag et al., 2021a). Then, Fomicheva et al. (2022) evaluate how well feature importance is correlated with human word-level error annotations, i.e., the plausibility of the explanations.
We note that, depending on the explainability method, these scores might be inverted, following the interpretation “errors are not important for a metric to achieve a high score”.111111Most gradient and attention based approaches return high scores for errors (e.g., Treviso et al., 2021; Rubino et al., 2021), for embedding-based approaches it depends on whether cosine-similarity or -distance is used (e.g., Leiter, 2021; Tao et al., 2022), for attribution based techniques using LIME (Ribeiro et al., 2016) or SHAP (Lundberg and Lee, 2017), low scores are returned for errors (e.g., as part of the baseline implementations in Eval4NLP21: https://github.com/eval4nlp/SharedTask2021/tree/main/baselines).